You bought a stock four months ago for $75.32 per share. The stock paid no dividends. The current share price is $78.34.

Required:
What is the APR and EAR of your investment? (Do not include the percent signs (%). Round your answers to 2 decimal places (e.g., 32.16).)


Investment
APR= %
EAR= %
Answer #1
The gain on this investment is $78.34 - 75.32 = $3.02
The percentage gain is 3.02/75.32 = 0.04009 = 4% for 6 months.
The APR is twice that, or 8%


The EAR is: (1+i/n)^n - 1 = (1+0.04)^2 - 1 = 0.0816 = 8.16%

so
APR= 8 %
EAR= 8.16 %
answered by: DeAndrea
Answer #2
The gain on this investment is

$78.34 - 75.32 = $3.02
The percentage gain is

3.02/75.32 = 0.0400 = 4% for 6 months
The APR is twice that, or 8%
The EAR is: (1+i/n)^n - 1
= (1+0.004)^2 - 1
= (1.0605)(1.0605) -1 = 0.008016 = 8.016%
answered by: eudora c
Answer #3

78.34 - 75.32 = $3.02
The percentage gain is

3.02/75.32 = 0.0400 = 4% for 4 months (which makes it quartly)
The APR is three times that, or 12%
The EAR is: (1+i/n)^n - 1
= (1+(0.12/3))^3 - 1
= (1.04)^3 -1 = 0.1249 = 12.49%